OpenStudy (anonymous):

solve x+3y=8 and 2x-y=2 using any method. show work.

OpenStudy (precal):

-2(x+3y=8) -2x-6y=-16 2x-y=2 -7y=-14 y=2 x+3y=8 x+3(2)=8 x+6=8 x=2 (2,2)

OpenStudy (anonymous):

thanks!!!! (:

OpenStudy (precal):

anytime :)

OpenStudy (anonymous):

how did you get the -2 at the beginning?